Find the dimensions of a rectangle whose width is 6 miles less than its length, and whose area is 91 square miles
Answer by mananth(16946)

Area = 91 sq. miles
Length = x miles
Width = x -6
Area = x * (x -6)
91 = x^2 -6x
x^2 -6 x -91 = 0
x^2-13x+7x-91=0
x(x-13)+7(x-13)=0
(x-13)(x+7)=0
x= 13 OR -7
Ignore negative value
Length =13 & width = 7miles